OpenStudy (anonymous):

Solve: x over 2 = 10 A) 20 B) 5 C) 1/2 D) 2

OpenStudy (anonymous):

(x/2) = 10 multiply both sides by 2 to isolate x what will you get ?

OpenStudy (anonymous):

is it 5?

OpenStudy (anonymous):

no (x/2) * 2 = 10 * 2 how the equation will look like ?

OpenStudy (anonymous):

so its 20

OpenStudy (anonymous):

yes ..

OpenStudy (anonymous):

thank you :)